A typical hot-air balloon has a volume of 15.8 m3 when inflated. How much work (in J) must be done to inflate the balloon when atmospheric pressure is 782 torr?

J
1
Expert's answer
2016-11-27T13:59:13-0500
Work is a result of multiplication force for distance.
P = 133.322368 * 782 Torr= 104258 Pa
1 Pa = 1 N/m2
W = 104258 * 15.8 = 1647276 Nm (J)